# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v3

@@ 3.3 Self-Training
To filter high-quality reasoning data from MCTS, we adapt two approaches: using the MC-Score or leveraging PRM/HRM to assign scores. To mitigate reward hacking caused by the reward model, we apply a high-quality data filter based on MC-Score.
Due to computational constraints, we do not employ RL methods such as PPO Schulman et al. (2017) or GRPO Shao et al. (2024). Instead, we continue using supervised fine-tuning. To preserve the general capabilities of the policy model, we incorporate causal language modeling loss combined with KL divergence regularization using a reference model. The objective function is defined as:
@@
Without proper KL regularization or with an insufficiently weighted KL loss (i.e., a very small $\lambda$), the KL divergence grows unbounded during training. Specifically, KL loss typically ranges from 00 to $20000$, whereas the causal LM loss remains within 00 to $12$, leading to a severe loss imbalance. This causes the optimization process to excessively minimize KL divergence at the expense of task-specific reasoning performance. To address this, we apply a logarithmic scaling to $D_{\text{KL}}(P||Q)$, stabilizing the loss landscape and ensuring a balanced trade-off between preserving general language capabilities and enhancing reasoning ability. Further details are provided in Section 4.3.
4 Experiment
4.1 HRM
Previous Step Label Current Step Label Label for Merged Step
Positive Positive Positive
Positive Neutral/Negative Negative
Neutral/Negative Positive Positive
Neutral/Negative Neutral/Negative Negative
Table 3: Accuracy of Qwen2.5-72B-Math-Instruct Yang et al. (2024b) evaluated under the best-of-$N$ strategy using ORM, PRM, and HRM on the PRM800K test set. ORM, PRM, and HRM are fine-tuned on the manually labeled PRM800K dataset.
@@ 4.1 HRM
Given that the PRM800K dataset Lightman et al. (2023) consists of Phase1 and Phase2, where Phase1 includes manually annotated reasoning processes, we utilize these manual annotations to construct the training datasets for ORM, PRM, and HRM. ORM training data comprises complete reasoning trajectories, while PRM training data consists of individual reasoning steps conditioned on preceding context. HRM training data extends PRM by incorporating multiple consecutive reasoning steps, allowing HRM to capture self-reflection and ensure reasoning coherence across sequential steps. Table 2 summarizes the labeling rules for merged reasoning steps in HRM.
We fine-tune Qwen2.5-1.5B-Math Yang et al. (2024b) as the reward model (RM) for classifying given reasoning step as correct or incorrect. Given an input, RM predicts logits for the positive and negative classes, denoted as $l_{\text{positive}}$ and $l_{\text{negative}}$, respectively. The confidence score is obtained by applying the softmax function:
$P(y=\text{positive}\mid x)=\frac{\exp(l_{\text{positive}})}{\exp(l_{\text{positive}})+\exp(l_{\text{negative}})},$ (2)
where $P(y=\text{positive}\mid x)$ denotes the probability of given reasoning step being correct. This probability serves as the reward assigned by RM. Detailed information is provided in Appendix A.1.
4.2 HNC
In this section, we utilize only the questions and ground truth from the PRM800K dataset Lightman et al. (2023), without relying on manually annotated data. We adopt MCTS as the automatic annotation method, using Qwen2.5-7B-Math-Instruct Yang et al. (2024b) as the reasoning engine to generate trajectories. As mentioned in Section 3.2, these auto-annotated reasoning trajectories from MCTS are used to train PRM, after which we apply the HNC data augmentation method to generate additional training data for HRM.
To balance computational efficiency and robustness, we configure MCTS with 5–6 child nodes per parent and a maximum tree depth of 7, ensuring reasoning completion within 7 steps. Since the computational cost of MCTS rollouts grows exponentially with tree depth and branching factor, we limit these parameters to maintain feasibility. The full MCTS simulation requires approximately 2,457 A100-80GB GPU-hours, while the HNC augmentation process takes around 30 minutes with CPU.
We perform supervised fine-tuning of Qwen2.5-1.5B-Math Yang et al. (2024b) for both PRM and HRM. To evaluate performance, we employ different policy models, including Qwen2.5-7B-Math-Instruct Yang et al. (2024b), DeepSeek-Math-7B Shao et al. (2024), and Qwen2.5-72B-Math-Instruct Yang et al. (2024b), applying the best-of-$N$ strategy on the PRM800K dataset. Detailed training information is provided in Appendix A.2.
Table 4: Accuracy of different policy models under PRM and HRM using the best-of-$N$ strategy on the PRM800K test set Lightman et al. (2023). The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct Yang et al. (2024b).
@@ 4.2 HNC
The relatively lower performance of Qwen2.5-72B-Math-Instruct (as shown in Table 4) can be attributed to the tree height constraints imposed by MCTS, which require answer generation to follow a predefined template within a fixed number of steps. While Qwen2.5-72B-Math-Instruct exhibits strong reasoning capabilities Yang et al. (2024b), its highly structured training process makes it more vulnerable to performance degradation when deviating from its learned format.
4.3 Self-Training
@@
We fine-tune Qwen2.5-7B-Math-Instruct Yang et al. (2024b) using high-MC-score reasoning data extracted from the PRM and HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on high-MC-score reasoning data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 5 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.
@@
4.4 HRM Generalization Across Different Domains
Table 6: Generalization performance of PRM and HRM fine-tuned on the PRM800K dataset and evaluated on GSM8K and MATH500 using the best-of-$N$ strategy. The policy model used for evaluation is Qwen2.5-7B-Math-Instruct.
@@ 5 Conclusion
In this paper, we present Hierarchical Reward Model, which enhances multi-step reasoning evaluation by integrating fine-grained and coarse-grained assessments, improving reasoning coherence and self-reflection. We further introduce Hierarchical Node Compression, a data augmentation method that optimizes MCTS-based autonomous annotation, enhancing label diversity while expanding training data with minimal computational cost. Extensive experiments on PRM800K dataset demonstrate HRM’s superior robustness over PRM, with strong generalization across GSM8K and MATH500 dataset. Additionally, MCTS-generated auto-labeled data enables policy model fine-tuning, further improving reasoning performance.
@@ Appendix A Appendix
A.1 HRM Training Details
To accelerate the training process of reward model, we employ FlashAttention Dao et al. [2022], Dao [2024], DeepSpeed Rajbhandari et al. [2020], Aminabadi et al. [2022], and mixed-precision training Kalamkar et al. [2019], Micikevicius et al. [2017]. However, within the PRM800K domain, we frequently encounter the issue: "Current loss scale already at minimum - cannot decrease scale anymore. Exiting run." This indicates that the numerical precision is insufficient for stable training. To mitigate this issue and ensure reproducibility, we set max_grad_norm to 0.01, which effectively stabilizes the training process.
@@ A.1 HRM Training Details
You are an expert of Math and need to solve the following question
and return the answer.
@@
Let’s analyze this step by step.
After you finish thinking, you need to output the answer again!
The answer should start with ’# Answer’, followed by two line breaks
and the final response.
Just provide the answer value without any descriptive text at the end.
@@
Below is a correct example of the expected output format:
@@ A.2 HNC Setting Details
"""You are an expert of Math and need to solve the following question
and return the answer.
@@
Let’s analyze this step by step.
Begin each step with ’# Step X’ to clearly indicate the entire reasoning step.
After you finish thinking, you need to output the answer again!
The answer should start with ’# Answer’, followed by two line breaks
and the final response.
Just provide the answer value without any descriptive text at the end.
@@
Below is a correct example of the expected output format:

### v4

@@ 3.3 Self-Training
We filter high-quality reasoning data from MCTS using the MC-Score to ensure reliable supervision and avoid reward bias. Due to computational constraints, we do not employ RL methods such as PPO Schulman et al. (2017) or GRPO Shao et al. (2024). Instead, we continue using supervised fine-tuning. To preserve the general capabilities of the policy model, we incorporate causal language modeling loss combined with KL divergence regularization using a reference model. The objective function is defined as:
@@
Table 2: Accuracy of Qwen2.5-72B-Math-Instruct on the PRM800K test set under the best-of-$N$ strategy, comparing ORM, PRM, and HRM. All models are fine-tuned on the manually labeled PRM800K dataset.
@@
Table 3: Accuracy of different policy models under PRM and HRM using the best-of-$N$ strategy on the PRM800K test set. The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct.
Without proper KL regularization or with an insufficiently weighted KL loss (i.e., a very small $\lambda$), the KL divergence grows unbounded during training. Specifically, KL loss typically ranges from 0 to $20000$, whereas the causal LM loss remains within 0 to $12$, leading to a severe loss imbalance. This causes the optimization process to excessively minimize KL divergence at the expense of task-specific reasoning performance. To address this, we apply a logarithmic scaling to $D_{\text{KL}}(P||Q)$, stabilizing the loss landscape and ensuring a balanced trade-off between preserving general language capabilities and enhancing reasoning ability. Further details are provided in Section 4.3.
4 Experiment
4.1 HRM
Given that the PRM800K dataset consists of Phase1 and Phase2, where Phase1 includes manually annotated reasoning processes, we utilize these manual annotations to construct the training datasets for ORM, PRM, and HRM. ORM training data comprises complete reasoning trajectories, while PRM training data consists of individual reasoning steps conditioned on preceding context. HRM training data extends PRM by incorporating multiple consecutive reasoning steps, allowing HRM to capture self-reflection and ensure reasoning coherence across sequential steps. Table 6 summarizes the labeling rules for merged reasoning steps in HRM.
We fine-tune Qwen2.5-1.5B-Math Yang et al. (2024b) as the RM for classifying given reasoning step as correct or incorrect. Given an input, RM predicts logits for the positive and negative classes, denoted as $l_{\text{pos}}$ and $l_{\text{neg}}$, respectively. The score is obtained by applying the softmax function:
$P(y=\text{pos}\mid x)=\frac{\exp(l_{\text{pos}})}{\exp(l_{\text{pos}})+\exp(l_{\text{neg}})},$ (4)
where $P(y=\text{pos}\mid x)$ denotes the probability of given reasoning step being correct. This probability serves as the reward assigned by RM. Detailed information is provided in Appendix A.2.
Statistical robustness.
The evaluation itself already provides extensive statistical averaging. Each dataset contains thousands of test problems (1k in GSM8K, 500 in MATH500, and over 800k annotated reasoning steps in PRM800K), which smooths random variation across samples. Notably, MATH500 and PRM800K consist of challenging high-school and university-level problems (e.g., calculus, linear algebra), whose solutions require multi-step reasoning rather than surface-level pattern matching or guessing. Moreover, in the Best-of-$N$ evaluation (up to $N=512$), each reasoning step expands into $N$ candidate paths, forming an exponentially large search tree. This hierarchical exploration implicitly averages over a vast number of trajectories, effectively functioning as multi-seed sampling and significantly reducing evaluation variance.
@@ 4.1 HRM
4.2 HNC
In this section, we utilize only the questions and ground truth from the PRM800K dataset, without relying on manually annotated data. We adopt MCTS as the automatic annotation method, using Qwen2.5-7B-Math-Instruct as the reasoning engine to generate trajectories. As mentioned in Section 3.2, these auto-annotated reasoning trajectories from MCTS are used to train PRM, after which we apply the HNC data augmentation method to generate additional training data for HRM.
To balance computational efficiency and robustness, we configure MCTS with 5–6 child nodes per parent and a maximum tree depth of 7, ensuring reasoning completion within 7 steps. Since the computational cost of MCTS rollouts grows exponentially with tree depth and branching factor, we limit these parameters to maintain feasibility. The full MCTS simulation requires approximately 2,457 A100 GPU-hours, while the HNC augmentation process takes around 30 minutes with CPU.
We perform supervised fine-tuning of Qwen2.5-1.5B-Math for both PRM and HRM. Training HRM adds roughly 30% more samples compared to PRM, increasing total training time from about 84 GPU hours to 120 GPU hours on A100 GPUs. This additional cost is negligible relative to the over 2400 GPU hours required for generating the reasoning trajectories via MCTS, which constitute the true computational bottleneck.
To evaluate performance, we employ different policy models, including Qwen2.5-7B-Math-Instruct Yang et al. (2024b), DeepSeek-Math-7B Shao et al. (2024), and Qwen2.5-72B-Math-Instruct Yang et al. (2024b), applying the best-of-$N$ strategy on the PRM800K dataset. Detailed training information is provided in Appendix A.3.
@@ 4.2 HNC
Table 5: Generalization performance of PRM and HRM trained on PRM800K and evaluated on GSM8K and MATH500 under the Best-of-$N$ setting using Qwen2.5-7B-Math-Instruct.
The relatively lower performance of Qwen2.5-72B-Math-Instruct (as shown in Table 3) can be attributed to the tree height constraints imposed by MCTS, which limit answer generation to a predefined template of at most 6 reasoning steps and require explicit output of "# Step X". While Qwen2.5-72B-Math-Instruct exhibits strong reasoning capabilities, its highly structured training process makes it more likely to generate outputs that deviate from the required format. As a result, some outputs cannot be retrieved using regex-based post-processing, thereby affecting the overall measured performance.
4.3 Self-Training
We fine-tune Qwen2.5-7B-Math-Instruct using reasoning data with MC-score >0.9>0.9, extracted from the PRM and HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on such data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 4 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.
4.4 HRM Generalization Across Different Domains
@@ 5 Conclusion
In this paper, we present HRM, which enhances multi-step reasoning evaluation by integrating fine-grained and coarse-grained assessments, improving reasoning coherence and self-reflection. We further introduce HNC, a data augmentation method that optimizes MCTS-based autonomous annotation, enhancing label diversity while expanding training data with minimal computational cost. Extensive experiments on PRM800K dataset demonstrate HRM’s superior robustness over PRM, with strong generalization across GSM8K and MATH500 dataset. Additionally, MCTS-generated auto-labeled data enables policy model fine-tuning, further improving reasoning performance.
6 Limitations
6.1 Merged Steps
The reason why we only merge two consecutive steps at a time is to maintain the simplicity and clarity of the labeling strategy. When merging more than two steps—such as combining four reasoning steps—the number of possible label combinations increases rapidly (e.g., one positive, two negative, one neutral), making it difficult to define unified labeling rules and leading to potential conflicts. This stands in sharp contrast to the straightforward rules shown in Table 6, where merging only two steps allows for clear and consistent label definitions.
@@ Appendix A Appendix
A.1 Labeling Rules for HRM Training Data
Table 6 summarizes the labeling rules for merging reasoning steps in HRM, as applied to the manually labeled PRM800K dataset.
Previous Step Label Current Step Label Label for Merged Step
Positive Positive Positive
Positive Neutral/Negative Negative
Neutral/Negative Positive Positive
Neutral/Negative Neutral/Negative Negative
A.2 HRM Training Details
To accelerate the training process of reward model, we employ FlashAttention Dao et al. (2022); Dao (2024), DeepSpeed Rajbhandari et al. (2020); Aminabadi et al. (2022), and mixed-precision training Kalamkar et al. (2019); Micikevicius et al. (2017). However, within the PRM800K domain, we frequently encounter the issue: "Current loss scale already at minimum - cannot decrease scale anymore. Exiting run." This indicates that the numerical precision is insufficient for stable training. To mitigate this issue and ensure reproducibility, we set max_grad_norm to 0.01, which effectively stabilizes the training process.
@@ A.2 HRM Training Details
You are an expert of Math and need to
solve the following question and return
the answer.
@@
Lets analyze this step by step.
After you finish thinking, you need to
output the answer again!
The answer should start with ’# Answer’,
followed by two line breaks and the
final response.
Just provide the answer value without
any descriptive text at the end.
@@
Below is a correct example of the
expected output format:
@@ A.3 HNC Setting Details
"""You are an expert of Math and need to
solve the following question and return
the answer.
@@
Lets analyze this step by step.
Begin each step with ’# Step X to
clearly indicate the entire reasoning
step.
After you finish thinking, you need to
output the answer again!
The answer should start with ’# Answer’,
followed by two line breaks and the
final response.
Just provide the answer value without
any descriptive text at the end.
@@
Below is a correct example of the
expected output format: